Student Evaluation

K. Subramani LDCSEE. West Virginia University, Morgantown, WV ksmani@csee.wvu.edu

1 Instructor’s comments

We have completed a course on the fundamentals of Algorithm design. Your feedback on the benefits of this course will play a vital role in determining how this course is offered again. It is therefore important that you fill in the rest of this form in as much detail as you can!

2 Review

1. Do you like the content of the course?

2. Is the course being covered at the correct pace?

3. Is the course too mathematical?

4. Are you satisfied with the instructor’s help?

5. Do you like the teaching style of the instructor?

6. Was the course organized to meet your expectations?

7. Would you appreciate similar courses being offered by by this instructor?

8. What are the features about this course that you would like to retain?

9. What are the feature about this course that you would like to change?

10. What are the attributes of the instructor that you would like to see more of?

11. What are the attributes that you would like changed?